Sarah10:

A company makes a profit of $50 per software program and $35 per video game. The company can produce at most 200 software programs and at most 300 video games per week. Total production cannot exceed 425 items per week. How many items of each kind should be produced per week in order to maximize the profit? Use linear programming to solve. Show all your work.

Shadow:

Computer science class?

Shadow:

Never mind, just haven't heard of this concept.

Sarah10:

algebra 2

Shadow:

Well you should start by giving each term a variable. Let x = the # of software programs Let y = the # of video games x is less than or equal to 200 y is less than or equal to 300 x + y is less than or equal to 425 Since x makes more $ you want as many of those as possible. Then it's all about fitting in the rest with y.

Sarah10:

how does x make more $ than y when its equal to or less than 200 ?

Shadow:

x, software programs, have a profit of $50 y, video games, have a profit of $35

Sarah10:

so this is what i only know x<= 200 y<= 300 x + y <= 425 but i dont know where to go from here .. how do i fit in the y? do i input them in the 50x ; 35y like this ? 50<=200 35<=300

Sarah10:

confusion consumes me 0_o

Shadow:

\[200 + y \le 425\] \[y \le 225\] 200 of x per week, 225 of y per week Both are within the constraints

Shadow:

You max profits by doing as many of the $50 as you can, then doing the rest as $35 ones

Sarah10:

do as many 50's without going over the limit?

Shadow:

Yes

Shadow:

I am sure there is a more eloquent way of approaching that @Vocaloid may know of but I'm kind of hungry and there is steak downstairs.

Shadow:

That's the general idea

Sarah10:

omg im sorry food always comes first thank you for the help i understand now!

Shadow:

It's all good. Glad I could help (: